OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 4 points) [2.05] MA.8.A.4.1 Tobias and Brian plan to share the hot chocolate from a full cylindrical thermos equally. The thermos has an inner radius of 1.5 inches and an inner height of 10 inches. If they will each use a mug with an inner radius of 2 inches, how high will the hot chocolate be in each mug to the nearest hundredth of an inch? Use the formula: V = 3.14r2h. 2.81 inches 2.97 inches 4.74 inches 5.62 inches

OpenStudy (souvik):

first calculate the volume of hot chocolate in thermos.. V=3.14*r^2*h

OpenStudy (anonymous):

and then solve the above for h

OpenStudy (anonymous):

V=3.14(1.5)^2(10) V=70.65 divide by two because they share the hot chocolate equally 70.65/2=35.325 subsitute in the original equation and solve for h(height) 35.325=3.14(2)^2h 35.325=12.56h 35.325/12.56=h h=2.81 Option A :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome :)